# Vizing’s Theorem in Near-Linear Time

### v1

@@ Appendix A Logarithmic Factors and Small Values of $\Delta$
For this step, we use the following result from [BD23] (itself based on [Ber22, Chr23]) that allows for coloring a random uncolored edge in essentially (but not exactly) $\operatorname*{poly}(\Delta)$ time. We note that a recent work of [BD24] presents a further improvement to these bounds (in terms of the exponent of $\Delta$) but as our dependence on $\Delta$ is only logarithmic, this improvement is inconsequential and thus we stick with the simpler work of [BD23].
Proposition A.7 ([BD23]).
@@
Lemma A.8.
Let $c$ be the $\operatorname*{poly}(\Delta)$-exponent in Proposition A.7, and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$. There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and $\lambda=\Theta(\lambda_{0})$ uncolored edges, extends the coloring to all uncolored edges in $O(m\log\Delta)$ time with high probability.
@@
Now consider the case when $\lambda_{0}=m/\Delta^{2c}$. We run the algorithm of Proposition A.7 with the following modification: when coloring a random edge, if the runtime becomes two times more than the expected time guarantee of Proposition A.7, we terminate the coloring and repeat from the beginning. Thus, each step takes $O(\Delta^{c}\cdot\log{(n/i)})$ time deterministically and succeeds in extending the coloring by one edge with probability at least half.
@@
While $X_{i}$’s are technically not independent (the choice of some edge being colored may change the structure of the graph for coloring next uncolored edge), they are stochastically dominated by independent random variables with above tail bounds (given the guarantee of Proposition A.7 for each uncolored edge). Thus, we can apply Proposition A.5 with parameters $\{\alpha_{i}\}_{i=1}^{\lambda}$ as above and $\beta_{i}=1$ for all $i\in[\lambda]$ to obtain that for $t:=m/\Delta^{c}$,
@@
Finally, we present an algorithm for taking care of the gap between the $m/\Delta$ uncolored edges left by Lemma A.6 and the $\Theta(1)\cdot\max\{m/\Delta^{2c},m\log{n}/n\}$ edges that Lemma A.8 can handle.
Lemma A.9.
Let $c=\Theta(1)$ and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$ be as in Lemma A.8. There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ with $\lambda=m/\Delta$ uncolored edges, extends the coloring to all but $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
@@
We simply run our Lemma A.4 with the given parameters $\lambda$ and $\lambda_{0}$. By Lemma A.4, the expected runtime will be some $T_{0}=O(m\log{(\lambda/\lambda_{0})}+\Delta\lambda)=O(m\log{\Delta})$ since $\lambda_{0}\geq m/\Delta^{2c}$ for some absolute constant $c\geq 1$. Moreover, by the same lemma,
Pr[runtime of the $\Pr\left[\text{runtime of the algorithm}\geq 2000\cdot T_{0}\right]\leq\exp\left(-10\lambda_{0}\right)\leq\exp\left(-\Theta(m\log{n})/n\right)=1/\operatorname*{poly}(n),$
Putting our above lemmas together concludes the proof of Theorem A.1 as follows:
Proof of Theorem A.1.
The proof follows immediately from the lemmas in Step 4:
  • By Lemma A.6, we can color all but $m/\Delta$ edges in $O(m\log{\Delta})$ time with high probability;
  • Then, by Lemma A.9, we can extend the coloring to all but $\Theta(\lambda_{0})$ edges of the graph in $O(m\log{\Delta})$ time with high probability, where $\lambda_{0}$ is as in Lemma A.8;
  • Finally, by Lemma A.8, we can color all the remaining $\Theta(\lambda_{0})$ uncolored edges in $O(m\log{\Delta})$ time with high probability.
This concludes the proof. ∎

### v3

@@ Appendix C Edge Coloring Algorithms for Small Values of $\Delta$
We now present the final part for coloring the last remaining edges. For this step, we use the following result from [BD23] (itself based on [Ber22, Chr23]) that allows for coloring a random uncolored edge in essentially (but not exactly) $\operatorname*{poly}(\Delta)$ time. We note that a recent work of [BD24] presents a further improvement to these bounds (in terms of the exponent of $\Delta$) but as our dependence on $\Delta$ is only logarithmic, this improvement is inconsequential and thus we stick with the simpler work of [BD23].
Proposition C.3 ([BD23]).
@@
Lemma C.4.
Let $c$ be the $\operatorname*{poly}(\Delta)$-exponent in Proposition C.3, and $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$. There is an algorithm that, given a graph $G$, a partial $(\Delta+1)$-edge coloring $\chi$ and $\lambda=\Theta(\lambda_{0})$ uncolored edges, extends the coloring to all uncolored edges in $O(m\log\Delta)$ time with high probability.
@@
Now consider the case when $\lambda_{0}=m/\Delta^{2c}$. We run the algorithm of Proposition C.3 with the following modification: when coloring a random edge, if the runtime becomes two times more than the expected time guarantee of Proposition C.3, we terminate the coloring and repeat from the beginning. Thus, each step takes $O(\Delta^{c}\cdot\log{(n/i)})$ time deterministically and succeeds in extending the coloring by one edge with probability at least half.
@@
While $X_{i}$’s are technically not independent (the choice of some edge being colored may change the structure of the graph for coloring next uncolored edge), they are stochastically dominated by independent random variables with above tail bounds (given the guarantee of Proposition C.3 for each uncolored edge). Thus, we can apply Proposition 10.3 with parameters $\{\alpha_{i}\}_{i=1}^{\lambda}$ as above and $\beta_{i}=1$ for all $i\in[\lambda]$ to obtain that for $t:=m/\Delta^{c}$,
@@
The proof of Corollary C.1 for $(\Delta+1)$ edge coloring is now as follows: run Lemma 10.4 to color all but $m/\Delta$ edges. Then, let $c$ be the constant in the exponent of $\operatorname*{poly}(\Delta)$ in Proposition C.3 and run Lemma C.2 to extend the coloring to all but $\lambda_{0}:=\max\{{m}/{\Delta^{2c}},m\log{n}/n\}$ edges. Finally, run Lemma C.4 to extend the coloring to all these remaining $\lambda_{0}$ edges. Since each of these algorithms take $O(m\mathrm{lo}{\Delta})$ time and succeed with high probability, we obtain the final algorithm.
C.2 Proof of Corollary C.1 for Multigraphs
To extend the proof to multigraphs, recall that by Lemma A.10 and Lemma B.10, we can, respectively, $(\Delta+\mu)$ color or $\left\lfloor 3\Delta/2\right\rfloor$ color all but $m/\Delta$ edges of a multigraph in $O(m\log{\Delta})$ time. Thus, we can follow the same exact strategy as in $(\Delta+1)$ coloring approach outlined in the previous subsection, and we only need an analogue of the result of [BD23] to conclude the proof. Such results are also already established by [Dha24a] and we do not need to reinvent the wheel here.
Proposition C.5 (A subroutine in [Dha24a, Theorem 1.6; part 1]).
There is an algorithm that, given a multigraph $G$, a partial $(\Delta+\mu)$-edge coloring $\chi$, and $\lambda$ uncolored edges, extends the coloring to a randomly chosen uncolored edge in expected $\operatorname*{poly}(\Delta)\cdot\log{(n/\lambda)}$ time.
Proposition C.6 (A subroutine in [Dha24a, Theorem 1.5; part 1]).
There is an algorithm that, given a multigraph $G$, a partial $\left\lfloor 3\Delta/2\right\rfloor$-edge coloring $\chi$, and $\lambda$ uncolored edges, extends the coloring to a randomly chosen uncolored edge in expected $\operatorname*{poly}(\Delta)\cdot\log{(n/\lambda)}$ time.
The rest of the proof is verbatim as in simple graphs in Corollary C.1 and we omit it here.
Appendix D A Specialized Concentration Inequality
Throughout the paper, for optimizing log factors in the runtime of our algorithms, we needed the following specialized concentration inequality. For completeness, we now prove this inequality using a standard argument based on moment generating functions.
Proposition (Restatement of Proposition 10.3).
then, for every $t\geq 0$,
@@ Appendix D A Specialized Concentration Inequality
for all $t\geq 0$. We further have,
$\mathbb{E}\left[e^{s\cdot Y_{i}}\right]$ $=\int_{y=0}^{\infty}\Pr\left[Y_{i}=y\cdot\alpha_{i}\right]\cdot\exp\left(s\cdot y\cdot\alpha_{i}\right)~dy$
$=\int_{y=0}^{\infty}\Pr\left[X_{i}=(1+y)\cdot\alpha_{i}\right]\cdot\exp\left(s\cdot y\cdot\alpha_{i}\right)~dy$ (by the connection between $X_{i}$ and $Y_{i}$ when $y\geq 0$)
$\leq\int_{y=0}^{\infty}\exp\left(-y\cdot\beta_{i}\right)\cdot\exp\left(s\cdot y\cdot\alpha_{i}\right)~dy$ (by the tail inequality of $X_{i}$ in the proposition statement)
$\leq\frac{1}{1-s\cdot(\alpha_{i}/\beta_{i})}$ (as $\beta_{i}\geq 1$)
Plugging this bounds in Equation 4, implies that
by the choice of $s$, concluding the proof. ∎